OpenStudy (anonymous):

Write the equation y=-1/2x + 3 in standard form

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

do you have any idea of where to start

OpenStudy (anonymous):

here is anotherone, Write the equation 9x +2y =6 in slope-intercept form

OpenStudy (anonymous):

did you get the first one

OpenStudy (anonymous):

no

OpenStudy (anonymous):

add -1/2x to both sides

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

so y1/2 = - 1/2x + 4.5

OpenStudy (anonymous):

close